So perpendicular lines have negative reciprocal slopes, but the 1st slope is (-1/4) so what would be the negative reciprocal?
Found 2 solutions by greenestamps, Alan3354:
Answer by greenestamps(13332)



Saying that the slope of a perpendicular line is the "negative reciprocal" of the slope of the first line does NOT mean the slope is a negative number. It means the slope of the perpendicular line is the reciprocal of the slope of the given line, MULTIPLIED BY NEGATIVE 1.

If the slope of the first line is -1/4, then the reciprocal of that slope is -4/1, and the slope of the perpendicular line is -1(-4/1) = 4/1 = 4.

Answer by Alan3354(69443)

Perpendicular lines have slopes with a product of -1
----
m1*m2 = -1 ------ (neither slope is zero)